2 Return on Investment The income from operations and the amount of invested assets in each...

2

Return on Investment

The income from operations and the amount of invested assets in each division of Beck Industries are as follows:

Income from Operations Invested Assets
Retail Division $129,200 $680,000
Commercial Division 56,000 350,000
Internet Division 151,800 660,000

a. Compute the return on investment for each division. (Round to the nearest whole number.)

Division Percent
Retail Division %
Commercial Division %
Internet Division %

b. Which division is the most profitable per dollar invested?
Answer #1

Return on investment = Income from operations / Invested assets

Retail division = 129,200/680,000 = 19%

Commercial division = 56,000/350,000 = 16%

Internet disvion = 151,800/660,000 = 23%

Internet disvion is more profitable